determine whether the graph of the two lines are parallel.
y+3=7x
3x-y=-3

Answer by mananth(16949)

y+3=7x
y=-7x-3...........1
3x-y=-3
y=3x+3
compare these equations with y=mx+b
the slopes are not equal.
so they are not parallel
m.ananth@hotmail.ca